If I already am given the solution of 3, -4, then how do I turn these numbers into a quadratic equation?
Answer by MRperkins(300)

Email me at justin.sheppard.tech@hotmail.com with the problem and I will reply with the work in a .pdf file. We can also set up a time to work on problems together online using a whiteboard that will allow you to work through the problem by yourself. All of this is free (volunteer).

Basically, you will have x-3=0 and x+4=0 so (x-3)(x+4)=0 FOIL and you will get your answer.
If you want more help, then I am happy to talk to you over the phone or via skype, both for free. Email with any questions or if you need to set up one-on-one help.
Justin